# RE-RAG: Improving Open-Domain QA Performance and Interpretability with Relevance Estimator in Retrieval-Augmented Generation

### v1

@@ 4.2 Effect of using less document
5 Analysis
5.1 Relationship between context relevance quality and answer confidence
We examine how the quality of the context set input to the generator relates to the confidence in the answer generation. We set the threshold for the generation probability of “true” tokens to 0.5, meaning that the generation probability of “true” tokens produced by the context RE is greater than that of “false” tokens. Then, we classify the input text as having high relevance quality if the context RE’s probability of generating a “true” token exceeds the threshold. For comparison to the baseline retriever, we use the cosine similarity of the hidden representation between the question and context in the retriever.
Figure 3 shows how the accuracy of answers varies with the proportion of high relevance quality among the top-25 input contexts. In both datasets, accuracy is increasing with the percentage of high relevance quality as measured by the context RE. In particular, accuracy decreases significantly in the absence of high relevance quality context. For the baseline retriever, we notice different behavior in two datasets. The baseline shows higher-than-expected performance for contexts with low relevance in TQA and lower-than-expected accuracy for high relevance contexts. This shows that the context RE can estimate the reliability of the final answer that will be generated by the system by measuring the context RE for the retrieved context in advance while baseline cannot.
5.2 Effectiveness of the context RE
We perform a qualitative analysis to see if our proposed context relevance estimator (RE) is effectively classifying relevant contexts. Table 3 shows a few contexts in the NQ test set.
Some of the contexts that the context RE predicts are highly relevant to the question even when they do not contain the exact ground truth answer. The first few examples in Table 3 are examples that are categorized as true context because they contain phrases that are semantically equivalent to the correct answer albeit not having the exact same form in the context. This shows that although the context RE is trained to measure the relevance of a question to a context through a limited set of ground truth answers, it is actually capable of measuring a broader range of relevance.
In addition to the examples above, there are cases where the context RE misclassified contexts as containing the correct answer. As shown in the example in Table 3, the context RE classified the context containing “the number of classes of strong verbs in German” as the correct context for the question about “the number of strong verbs in German”, which means that our context RE is still limited in its ability to capture the fine-grained meaning of the question in the retrieved context. On the other hand, in the last example, for the question about “the number of episodes”, it succeeded in classifying the context containing “the number of classical episodes” as an incorrect context.
5.3 RE for classifying “irrelevant” context set
Table 4 shows the performance of the context relevance estimator (RE) as a “irrelevant” set classifier. “irrelevant” set means that the context set of the top-25 contexts input to the generator does not contain an answer in any context. For classification, we used the cosine similarity score of the hidden representation of the question and context for retriever and the probability of generating a "true" token by the model for context RE. For the optimal threshold, we searched for the value that maximizes F1 score in steps of 0.1 from 0.5 to 0.9.
Our context RE showed better “irrelevant” set classification performance than FiD-KD’s improved retriever based on F1 score. Looking at the detailed performance, we found that the retriever performed better for recall, but the context RE performed better for precision. This is because the retriever classified a large number of context sets as all “irrelevant” sets, while our proposed context RE showed a good balance between classification precision and recall.
Table 5 shows the accuracy changes after making the model respond with “unanswerable” to “irrelevant” sets, divided into context sets with and without answers. Model set to respond “unanswerable” if no context exceeds threshold set in Table 4. The accuracy in sets where answers can be found slightly decreases due to incorrect “unanswerable” responses. Conversely, in sets where answers cannot be found, accuracy increases (from 0) by responding with “unanswerable” to “irrelevant” sets, thereby improving accuracy in cases where answers are unattainable.
| Dataset | Model | Recall | Precision | F1 |
| NQ | FiD-KD | 73.2 | 21.9 | 33.7 |
| 51.3 | 33.9 | 40.9 | ||
| 45.9 | 38.3 | 41.7 | ||
| TQA | FiD-KD | 64.3 | 24.5 | 35.5 |
| 38.9 | 46.7 | 42.5 | ||
| 39.0 | 43.2 | 41.0 |
| Dataset | Model | relevant context set | |
| O | X | ||
| NQ | 58.3 54.9 | 51.3 | |
| 61.5 57.9 | 45.9 | ||
| TQA | 78.7 77.0 | 38.9 | |
| 80.4 77.9 | 39.0 | ||
5.4 Plugging the context RE into the LLM
We investigate whether the context relevance estimator (RE), which has been effective in relatively small-sized models, is also effective in improving LLM’s performance. We follow the method proposed in REPLUG (Shi et al., 2023) to marginalize the answers generated by the LLM according to the input contexts. We use OpenAI "gpt-3.5-turbo-0125" (Brown et al., 2020), which generates answers using the top- contexts evaluated by our context RE and evaluated by FiD-KD’s retriever. The score for each context was calculated using the "true" token logit from the context RE and the cosine similarity of the hidden representation of the question and context generated by the retriever. We used 8-shot prompts for NQ and 2-shot prompts for TQA. The detailed prompts are shown in Appendix C.
In both datasets, our proposed context RE outperforms the baseline retriever on both datasets. The difference in EM score was 6.4 for NQ and 2.8 for TQA. This indicates that context RE’s improved ability to re-rank contexts and its ability to calculate more accurate relevance scores can improve overall answer quality by simply combining context RE with LLM’s RAG system.
| Model | top-5 | top-10 | ||
| NQ | TQA | NQ | TQA | |
| GPT w/ Retriever | 41.7 | 67.3 | 42.9 | 69.0 |
| GPT w/ RE | 48.8 | 70.7 | 49.3 | 71.8 |
5.5 Ablation Study
We perform an ablation study to investigate the effectiveness of the added context RE in RE-RAG. The effect of our proposed context RE is twofold. First, it performs better re-ranking than the retriever, selecting more accurate context and passing it to the generator. Second, it calculates a more accurate relevance score than retriever’s similarity score and uses it in the answer marginalization process. In Table 7, the performance of methods with each component of the context RE added is presented, using a model that was trained with only the T5-base generator, after removing the context RE, as the baseline.
We construct the following experiment to isolate the two effects. First, we apply the top 25 contexts from retriever and their similarity scores to the baseline model. Next, there are the top-25 contexts from the retriever with the context RE’s score applied (RE score) and the top-25 contexts from the context RE with the retriever’s similarity score applied (RE context). Finally, we compare the performance of applying the context RE’s top-25 contexts and score to the baseline model (RE context, score).
Both effects of the context RE are found to be significant in improving the performance of the final model. This shows that not only the quality of the context input to the generator plays an important role, but also the score, which means the importance of each context. The effect of context RE is 7.3 EM for NQ and 9.0 EM for TQA. The impact of Context score is 3.6 EM for NQ and 5.2 EM for TQA.
| Model | NQ | TQA |
| Baseline | 39.5 | 54.9 |
| Baseline w/ RE score | 43.1 | 60.1 |
| Baseline w/ RE context | 46.8 | 63.9 |
| Baseline w/ RE context, score | 49.6 | 67.8 |
| 49.9 | 68.2 |
6 Related Works
Previous research has shown that the performance of Question Answering systems can be improved by utilizing external knowledge about questions (Chen et al., 2017). Methods for more accurate retrieval of external knowledge (Karpukhin et al. (2020); Khattab et al. (2021); Gao and Callan (2022)) have been studied to make these systems more efficient. In open-domain QA, models that extract and use answers from retrieved documents have been studied (Karpukhin et al. (2020); Khattab et al. (2021); Cheng et al. (2021)), but studies that utilize generative models such as T5 (Raffel et al., 2020) or BART (Lewis et al., 2020a) have become more common (Lewis et al. (2020b); Izacard and Grave (2021b)). RAG and FiD achieved state-of-the-art performance in open-domain QA using different methods. Subsequently, models (Izacard and Grave (2021a); Jiang et al. (2022); Fajcik et al. (2021)) that leverage and improve upon the structural advantages of FiD have been proposed. For Atlas (Izacard et al., 2022), state-of-the-art performance was achieved through an improved retriever (Izacard et al., 2021) and scaling up the model. In the case of RAG, there is a study that improved performance by introducing a BERT (Devlin et al., 2019)-based reranker (Glass et al., 2022), but it utilized additional data and high-quality label data when training the reranker. Recently, LLMs such as GPT (Brown et al., 2020) and Llama2 (Touvron et al., 2023), which have been developed in recent years, face limitations with FiD methods that require encoded data. Consequently, research on RAG models, which can directly input context, has received renewed attention. (Shi et al. (2023); Asai et al. (2023); Lin et al. (2023)) These approaches have achieved performance improvements by training a retriever, which can also be applied to LLM, or by performing the review of questions and context within the model itself.
7 Conclusion
We propose RE-RAG, which adds a context RE to the RAG system to evaluate the relevance between question and context. We show that our proposed RE-RAG achieves competitive performance on NQ and TQA datasets, and that the context RE can be combined into LLM independently to improve performance. Furthermore, RE-RAG is relatively easy to train as it does not utilize label data such as whether the context contains an answer or not to train the context RE.
8 Limitation
Our research has explored how to improve performance and interpretability by evaluating the relevance of questions and context in a original RAG model. Our work has focused on RAG systems, with limited exploration of how our methods can improve FiD-based systems. We believe that such research could be conducted in the future.
Appendix A Dataset Statistics
| Dataset | Train | Dev | Test |
| Natural Questions | 79,168 | 8,757 | 3,610 |
| TriviaQA | 78,785 | 8,837 | 11,313 |
@@ Appendix B Training Details
For model selection, we evaluated every 1 epoch and selected the case with the highest answer accuracy of the dev set. The dev set answer accuracy was measured using the top-10 context of the context RE. Since the answer accuracy of the top-10 context of the context RE is similar to the answer accuracy of the top-25 context, this helped to save computational resources and time while still producing valid results.
Appendix C Prompts utilized in GPT-3.5
Table 9 and Table 10 show the prompts provided in GPT-3.5 used in our experiments. We provided an 8-shot example for Natural Questions and a 2-shot example for TriviaQA. We performed a simple normalize on the sequence generated by the GPT model with the following prompts, and then treated as correct the answer that exactly matched the answer in the dataset.
Appendix D Relationship between context relevance quality and answer confidence at base model
Figure 3 illustrates the relationship between context relevance quality and answer confidence described in Section 5.1 in . The overall trend is not significantly different from the large model, but we can see that the difference in accuracy is caused by the difference in the number of parameters in the generator model.
| ###Description : Below are some examples of question and answer formats. Use these examples as a guide to help you come up with the right answer to the question you’ll eventually be asked. |
| Example 1 |
| Context: ’title’: Service club, ’text’: "the services", a common expression for the military or uniformed forces. In the Americas, these types of clubs are commonly known as veteransórganizations or veteransf́raternal groups. The worldś first service club, the Rotary Club of Chicago, was formed in 1905 by Paul P. Harris, an attorney who wanted to create in a professional club with the same friendly spirit he had felt in the small towns of his youth. The Rotary name derived from the early practice of rotating meetings among membersóffices. Many of these service clubs were started early in the 20th century, such as Kiwanis, |
| Question: In which city were Rotary Clubs set up in 1905? |
| Answer: Chicago |
| Example 2 |
| Context: ’title’: Jason Schwartzman, ’text’: he played a writer who moonlights as an unlicensed private detective by advertising himself on Craigslist. He currently releases music through his solo project Coconut Records, and was formerly the drummer of rock band Phantom Planet. Schwartzman was born in Los Angeles, California, the son of actress Talia Shire (née Coppola) and the late producer Jack Schwartzman. Schwartzman’s brother is actor and musician Robert Schwartzman, and his paternal half-siblings are Stephanie and cinematographer John Schwartzman. Many other members of Schwartzman’s family are involved in film: he is the nephew of Francis Ford Coppola, cousin of Nicolas Cage, Sofia Coppola, Roman |
| Question: Which famous brother of Talia Shire does not share her last name? |
| Answer: Francis Ford Coppola |
| ###Instructions: Provide the correct answer to the given question. The given question is accompanied by context related to the question. Be sure to refer to the context provided and enter your answer based on the context after "###Answer:". The question and the answer you provide must be relevant. Write your answer in a short "short answer". |
| ###Context: {context} |
| ###Question: {question} |
| ###Answer: |
| ###Description : Below are some examples of question and answer formats. Use these examples as a guide to help you come up with the right answer to the question you’ll eventually be asked. |
| - Example 1 |
| Context: ’title’: Sports in the United States, ’text’: Erving (won MVP awards in both the ABA and NBA), Kareem Abdul-Jabbar (6 time MVP), Magic Johnson (3 time MVP), Larry Bird (3 time MVP), Michael Jordan (6 time finals MVP), John Stockton (#1 in career assists and steals), Karl Malone (14 time all NBA team), Kobe Bryant (NBA’s third all-time leading scorer), Tim Duncan (15-time NBA all-star), Shaquille O’Neal (3 time finals MVP) and Jason Kidd (#2 in career assists and steals). Notable players in the NBA today include LeBron James (4 MVP awards), Stephen Curry (2 time MVP), Dwyane Wade (10 time all-star), and Kevin Durant (MVP, 4 |
| Question: who are the top 5 leading scorers in nba histor |
| Answer: Kobe Bryant |
| Example 2 |
| Context: ’title’: My Hero Academia: Two Heroes, ’text’: would be joining the cast as Melissa Shield and Katsuhisa Namase would play David Shield, both original characters. On June 11, 2018, "Weekly Shōnen Jump" announced that Rikiya Koyama had been cast as the filmś villain, Wolfram. Masaki Suda performs the filmś theme song , which was written and composed by Hiromu Akita of amazarashi. Funimation and Toho premiered the film at Anime Expo in Los Angeles on July 5, 2018, and it was later released in Japan on August 3 of that year. The first one million audience members to see the movie will receive a special book containing |
| Question: when does the new my hero academia movie come out |
| Answer: July 5, 2018 |
| Example 8 |
| Context: ’title’: King Kong, ’text’: and the subsequent appeal. Since the court case, Universal still retains the majority of the character rights. In 1986 they opened a King Kong ride called "King Kong Encounter" at their Universal Studios Tour theme park in Hollywood (which was destroyed in 2008 by a backlot fire), and followed it up with the Kongfrontation ride at their Orlando park in 1990 (which was closed down in 2002 due to maintenance issues). They also finally made a King Kong film of their own, "King Kong" (2005). In the summer of 2010, Universal opened a new 3D King Kong ride called at |
| Question: when did the king kong ride burn down |
| Answer: 2008 |
| ###Instructions: Provide the correct answer to the given question. The given question is accompanied by context related to the question. Be sure to refer to the context provided and enter your answer based on the context after "###Answer:". The question and the answer you provide must be relevant. Write your answer in a short "short answer" of 5 words or less. |
| ###Context: {context} |
| ###Question: {question} |
| ###Answer: |

### v2

@@ Appendix B Training Details
For model selection, we evaluated every 1 epoch and selected the case with the highest answer accuracy of the dev set. The dev set answer accuracy was measured using the top-10 context of the RE. Since the answer accuracy of the top-10 context of the RE is similar to the answer accuracy of the top-25 context, this helped to save computational resources and time while still producing valid results.
| P-Generator | R-Generator | NQ | TQA |
| (N31.1/T64.3) | 36.1 35.8(-0.3) | 58.4 62.8(+4.4) | |
| 38.8 36.9(-1.9) | 64.9 65.4(+0.5) | ||
| 40.7 37.4(-3.3) | 66.3 66.2(-0.1) | ||
| (N41.3/T75.1) | 38.2 42.1(+3.9) | 57.6 66.9(+9.3) | |
| 46.8 45.6(-1.2) | 72.1 74.0(+1.9) | ||
| ChatGPT (N37.7/T72.0) | ChatGPT | 45.9 43.2(-2.7) | 70.7 72.1(+1.4) |
Appendix C Effectiveness of the RE
We perform a qualitative analysis to see if our proposed relevance estimator (RE) is effectively classifying relevant contexts. Table 3 shows a few contexts in the NQ test set.
Some of the contexts that the RE predicts are highly relevant to the question even when they do not contain the exact ground truth answer. The first few examples in Table 3 are examples that are categorized as true context because they contain phrases that are semantically equivalent to the correct answer albeit not having the exact same form in the context. This shows that although the RE is trained to measure the relevance of a question to a context through a limited set of ground truth answers, it is actually capable of measuring a broader range of relevance.
In addition to the examples above, there are cases where the RE misclassified contexts as containing the correct answer. As shown in the example in Table 10, the RE classified the context containing “the number of classes of strong verbs in German” as the correct context for the question about “the number of strong verbs in German”, which means that our RE is still limited in its ability to capture the fine-grained meaning of the question in the retrieved context. On the other hand, in the last example, for the question about “the number of episodes”, it succeeded in classifying the context containing “the number of classical episodes” as an incorrect context.
| Dataset | Type | Threshold | ||||
| 0.5 | 0.6 | 0.7 | 0.8 | 0.9 | ||
| NQ | Answerable | 61.3 | 56.2 | 34.9 | 6.4 | 0.0 |
| Unanswerable | 2.3 | 27.8 | 71.3 | 97.2 | 99.8 | |
| TQA | Answerable | 77.3 | 51.6 | 9.2 | 0.1 | 0.0 |
| Unanswerable | 14.3 | 62.7 | 94.7 | 100.0 | 100.0 | |
Appendix D Selectively using parametric knowledge with FiD-KD
Table 11 shows the change in EM score when applying the mixed decoding strategy, using the cosine similarity score of the FiD-KD retriever as the confidence score. For small parameter generators, the EM score is low when applying the FiD-KD retriever to LLMs, which results in a high gain when utilizing parametric knowledge of large parameter models. However, since the classification performance of the FiD-KD retriever is lower than that of RE, even utilizing parametric knowledge does not significantly outperform the baseline performance of parametric knowledge. Especially for more knowledge-intensive tasks such as NQ, the performance loss is substantial.
Appendix E FiD-KD retriever’s performance in “unanswerable” scenarios
Table 12 shows the performance of the FiD-KD retriever in unanswerable scenarios according to different threshold values. For the FiD-KD retriever, it is observed that while trying to maintain performance on the answerable set, the classification ability on the unanswerable set significantly decreases.